# Poisson Midpoint Method for Log Concave Sampling: Beyond the Strong Error Lower Bounds

### v2

@@ 1 Introduction
Sampling from a density $\pi(x)\propto\exp(-F(x))$ over $\mathbb{R}^{d}$ is of fundamental interest in physics, economics, and finance johannes2010mcmc ; von2011bayesian ; kobyzev2020normalizing . Applications in computer science include volume computation vempala2010recent and bandit optimization russo2018tutorial . A popular approach is Langevin Monte Carlo (LMC) which is the Euler-Maruyama discretization of the continuous time Ito Stochastic Differential Equation (SDE) called (overdamped/underdamped) Langevin Dynamics. The convergence of LMC has been extensively studied in the literature durmus2018analysislangevinmontecarlo ; vempala2022rapidconvergenceunadjustedlangevin ; erdogdu2021convergencelangevinmontecarlo ; cheng2017convergencelangevinmcmckldivergence ; cheng2018-underdamped-lmc ; dalalyan2020sampling ; altschuler2025shiftedcompositionivunderdamped under various assumptions on the potential function $F$ such as log concavity and isoperimetry.
@@
Strong $L^{2}$ error is the $L^{2}$ distance between the continuous time Ito SDE solution at time $T$ and the sampling algorithm output whenever they are driven by same Brownian motion. However, sampling algorithm guarantees generally consider ‘weak’ notions of distance such as total variation distance, Wasserstein distance, or the KL divergence between the law of algorithm output and the target. In particular, Wasserstein-2 distance bounds can consider the $L^{2}$ distance between algorithm output and the continuous time SDE driven by different but arbitrarily coupled Brownian motions.
@@ 2 Notation and problem setup
[…]
@@
We define the condition number $\kappa:=\frac{L}{\mu}$. Our notion of complexity is the number of gradient calls of $F$, in terms of the problem parameters $\kappa,d$ and $\epsilon.$ The remainder of this section closely follows the discussion in kandasamy2024poissonmidpointmethodlangevin , since we have the same problem setup.
2.1 Langevin Monte-Carlo
@@
Here $Z_{t}$ are independent standard Gaussians, $b(x,t):\mathbb{R}^{d}\times\mathbb{R}^{+}\to\mathbb{R}^{d}$ is a drift function; $A_{\eta}$ and $G_{\eta}$ are $d\times d$ matrices. We focus on two specific cases of the above.
Overdamped LMC: Here $A=\mathbf{I}_{d}$, $G_{\eta}=\eta\mathbf{I}_{d}$, $\Gamma_{\eta}=\sqrt{2\eta}\mathbf{I}_{d},$ and $b(x,t)=-\nabla F(x)$ for some potential $F.$ This is the Euler-Maruyama discretization of Overdamped Langevin dynamics:
@@
$dU_{t}=V_{t}dt,\text{ }dV_{t}=-\gamma V_{t}-\nabla F(U_{t})dt+\sqrt{2}dB_{t}$ (3)
The stationary distribution of these dynamics is $\pi(u,v)\propto\exp(-F(u)-\frac{1}{2}||v||^{2}).$ eberle2019couplings ; dalalyan2020sampling
@@ 2.2 The Poisson midpoint method
For $K\in\mathbb{N}$. Consider the dynamics in (2) with step size $\eta/K$, $X_{t}$. The Poisson midpoint method uses a randomized algorithm to approximate $K$ steps of the dynamics in with just 1 step of step size $\eta$ using 2 evaluations of the drift $b()$ in expectation. Denoting the iterates of the Poisson midpoint method as $\tilde{X}_{tK}$, this means $\tilde{X}_{tK}$ approximates $X_{tK}$. The work of kandasamy2024poissonmidpointmethodlangevin gives an efficient implementation of PLMC, with 2 evaluations of the drift $b()$ per $K$ steps to obtain $\tilde{X}_{(t+1)K}$ directly from $\tilde{X}_{tK}$. We consider the ‘fictitious’ interpolation $\hat{\tilde{X}}_{tK+i}$ (which approximates $X_{tK+i}$) to enable mathematical analysis. The iterates run in batches of size $K,$ denoted ${\tilde{X}}_{tK+i}$ with $0\leq i<K$. Let $Y_{tK+i}$ denote independent standard Gaussians. Given $\tilde{X}_{tK}$, the 0-th iterate of a batch, we define the cheap interpolation
@@
We consider this approximation in the two specific cases of Langevin Dynamics to obtain Poisson Langevin Monte-Carlo (PLMC).
Overdamped PLMC: Let $Z_{tK+i}$ be independent standard Gaussians, and $X_{tK+i}$ denote the iterates of Langevin monte-carlo, with step-size $\eta/K.$
@@
Underdamped PLMC: Let $Z_{tK+i}$ be independent standard Gaussians in $\mathbb{R}^{2d},$ and $X_{tK+i}=[U_{tK+i},V_{tK+i}]^{T}$ denote the iterates of underdamped Langevin Monte-Carlo with step-size $\eta/K$.
@@ 2.3 Prior work
LMC (durmus2018analysislangevinmontecarlo ) Log-Concave $W_{2}^{2}\leq\frac{\epsilon^{2}}{\alpha}$ $\frac{\kappa d}{\epsilon^{2}}$
RLMC (shen2019randomizedmidpointmethodlogconcave ; yu2023langevinmontecarlostrongly ) Log-Concave $W_{2}^{2}\leq\frac{\epsilon^{2}}{\alpha}$ $\frac{\kappa\sqrt{d}}{\epsilon}+\frac{\kappa^{4/3}d^{1/3}}{\epsilon^{2/3}}$
RLMC (altschuler2024shiftedcompositioniiilocal ) Log-Concave $\mathsf{KL}\leq\epsilon^{2}$ $\frac{\kappa\sqrt{d}}{\epsilon}$
@@
PLMC (Ours) Log-Concave $W_{2}^{2}\leq\frac{\epsilon^{2}}{\alpha}$ $\frac{\kappa^{4/3}d^{1/3}+\kappa d^{2/3}}{\epsilon^{2/3}}$
@@
LMC (dalalyan2020sampling ) Log-Concave $W_{2}^{2}\leq\frac{\epsilon^{2}}{\alpha}$ $\frac{\kappa^{3/2}\sqrt{d}}{\epsilon}$
RLMC (shen2019randomizedmidpointmethodlogconcave ; yu2023langevinmontecarlostrongly ) Log-Concave $W_{2}^{2}\leq\frac{\epsilon^{2}}{\alpha}$ $\frac{\kappa d^{1/3}}{\epsilon^{2/3}}+\frac{\kappa^{7/6}d^{1/6}}{\epsilon^{1/3}}$
@@
PLMC (Ours) Log-Concave $W_{2}^{2}\leq\frac{\epsilon^{2}}{\alpha}$ $\frac{\kappa^{7/6}d^{1/3}}{\epsilon^{1/3}}+\frac{\kappa^{\frac{11p+6}{8p+6}}d^{\frac{3p+2}{8p+6}}}{\epsilon^{\frac{p+2}{4p+3}}}$
Recent works have focused on the rigorous theoretical analysis of classical and popular MCMC algorithms to establish complexity bounds and theoretical limits. The prototypical case of Overdamped LMC has been studied when the target $\pi$ is strongly log-concave and more generally $\pi$ when satisfies isoperimetric inequalities dalalyan2017theoretical ; durmus2017unadjusted ; durmus2018analysislangevinmontecarlo ; vempala2022rapidconvergenceunadjustedlangevin ; erdogdu2021-tail-growth ; mou2022improved ; sinho-non-log-concave-lmc . Underdamped LMC has been considered as a faster alternative. This case too has been well studied when $\pi$ is strongly log-concave and when $\pi$ satisfies isoperimetric inequalities cheng2018-underdamped-lmc ; dalalyan2020sampling ; ganesh2020faster ; ma2021there ; zhang2023improveddiscretizationanalysisunderdamped ; altschuler2025shiftedcompositionivunderdamped
@@
Note that the literature on MCMC considers 3 notions of convergence: KL-divergence, TV distance and 2-Wasserstein distance. In the case when $\pi$ is strongly log-concave, the Otto-Villani Theorem otto2000generalization shows that $\mathsf{KL}(\mu||\pi)\leq\epsilon^{2}\implies W_{2}^{2}(\mu,\pi)\lesssim\frac{\epsilon^{2}}{\alpha}$ and the Pinsker’s inequality shows that $\mathsf{KL}(\mu||\pi)\leq\epsilon^{2}\implies\mathsf{TV}(\mu,\pi)\lesssim\epsilon$. The condition of $\pi$ satisfying LSI is more general than log-concavity of the target. We refer to Table 1 for a detailed comparison of the results.
@@ 2.4 Our contributions
However, sampling algorithm guarantees consider ‘weak errors’ which are distances between $\mathsf{Law}(U_{T}(\omega))$ and $\mathsf{Law}(A(F,\omega,\tilde{\omega}))$. In particular, the Wasserstein-2 distance is the infimum of $L^{2}$ errors when $U_{T}$ is driven by $B_{t}(\omega)$ and $A(\cdot)$ queries $B_{t}^{\prime}(\omega)$ over all couplings of distinct Brownian motions $B_{t}(\omega)$ and $B^{\prime}_{t}(\omega)$. Our results for Poisson ULMC shows queries the oracle $\tilde{O}_{\kappa,d}(\frac{1}{\epsilon^{1/3}})$ times in expectation to achieve $W_{2}^{2}(\mathsf{Law}(A(F,\omega,\tilde{\omega})),\pi)\leq\frac{\epsilon^{2}d}{\alpha}$, a quadratic improvement.
@@ 3 Results
Let $\tilde{X}_{tK+i}$ denote the iterates of Overdamped PLMC, and $X_{tK+i}$ the iterates of Overdamped LMC with stepsize $\eta/K.$ and $\mathsf{Law}(X_{0})=\mathsf{Law}(\tilde{X}_{0})$. Assume $\eta L\leq 1/8,$ and Assumption 1. Then there exist absolute constants $c_{1}$ and $c_{2}$ such that
@@
$\tfrac{\eta^{3}L^{4}d}{\alpha^{2}}+\tfrac{\eta^{4}L^{4}d^{2}}{\alpha}+\exp(c_{1}d-(c_{2}\eta^{2}L^{2}K)^{-1})\cdot\tfrac{\eta^{2}L^{2}d}{\alpha}.$
@@
Let $\tilde{X}_{t}$ be the iterates of Overdamped PLMC as in Theorem 1. Assume the conditions from Theorem 1, and that $0<\epsilon\lesssim\min(1,\kappa^{-1/4},d^{-1/4}).$ Let $x^{*}$ be the unique minimizer of $F$, and let $\tilde{X}_{0}$ satisfy $\mathbb{E}[F(\tilde{X}_{0})-F(x^{*})]\leq C_{F}\kappa d$ for some $C_{F}>0$. Then there exist constants $C_{1},C_{2}>0$ depending only on $C_{F}$, $\log(\frac{W_{2}(U_{0},\pi)\sqrt{\alpha}}{\epsilon\sqrt{d}})$ and $\log(1/\epsilon),$ polynomially, such that if $\eta=C_{1}\min(\frac{\alpha^{1/3}\epsilon^{2/3}}{L^{4/3}},\frac{\epsilon^{2/3}}{d^{1/3}L}),$ $K=4\frac{\eta L}{\epsilon^{2}}$ and $N=C_{2}\big{[}\frac{\kappa^{4/3}+\kappa d^{1/3}}{\epsilon^{2/3}}\big{]}$. Then,
@@
The complexity bound for Overdamped LMC (durmus2018analysislangevinmontecarlo ) is $\tilde{O}(\kappa/\epsilon^{2})$ gradient calls, and that of Overdamped RLMC (yu2023langevinmontecarlostrongly ) is $\tilde{O}(\frac{\kappa}{\epsilon}+\frac{\kappa^{4/3}}{\epsilon^{2/3}})$ gradient calls. To our knowledge, our method is thus the best known discretization of overdamped Langevin dynamics, in terms of $\epsilon$ dependence. Note that assumption our on the initialization is very mild - $F$ can be optimized quickly under Assumption 1.
@@
Let $\tilde{X}_{tK+i}$ denote the iterates of Underdamped PLMC, and $X_{tK+i}$ denote the iterates of Underdamped LMC with step-size $\eta/K$. Let $M=\begin{bmatrix}\mathbf{I}_{d}&0\\ \mathbf{I}_{d}&\frac{2}{\gamma}\mathbf{I}_{d}\end{bmatrix},$ and $p\geq 0$ be an integer. Assume $\eta/K\lesssim\frac{1}{\kappa\sqrt{L}},$ the conditions of Theorem 6, Assumption 1, and that $\gamma\eta<c_{0}$ for sufficiently small $c_{0}>0.$ Then with $\gamma=c\sqrt{L}$ for some $c\geq 2$, we have
@@
$+\mathbb{E}[P(||V_{0}||,|F(\Psi_{0})-F(\Psi_{t})|^{+})],$
Where $P$ is a polynomial, and $\Psi$ is defined as $\Psi_{s}:=\tilde{U}_{sK}+\frac{1}{\gamma}\tilde{V}_{sK}.$ The complete bound is laid out in Section 6.1, for the sake of clarity.
Similar to Corollary 1, the following Corollary (proved in Section F) establishes complexity bounds.
@@
Let ${\tilde{X}}_{tK+i}$ denote the iterates of Underdamped PLMC, as in Theorem 3. Assume the conditions of Theorem 3, and that $\epsilon\lesssim\min(1,\kappa d^{-1/2},\kappa^{\frac{p+3}{2(3p+1)}}d^{-\frac{p}{3p+1}})$. Set $K=\frac{\eta L\sqrt{2}}{0.94\epsilon\sqrt{\alpha}},$ $\gamma\asymp\sqrt{L}$, and initialize $V_{0}\sim\mathcal{N}(0,\mathbf{I}_{d}).$ Let $x^{*}$ be the unique minimizer of $F$, and initialize $X_{0}$ with $\mathbb{E}||X_{0}-x^{*}||^{2k}\lesssim d^{k}/L^{k}$ for $k=\max(2,p+1)$. Then there exist constants $C_{3},C_{4}>0$ depending only on $\log(\frac{W_{2}(U_{0},\pi)\sqrt{\alpha}}{\epsilon\sqrt{d}})$ and $\log(\frac{\kappa}{\epsilon}),$ polynomially, such that if $\eta=C_{3}\min\left(\frac{\epsilon^{1/3}}{\kappa^{1/6}d^{1/6}\sqrt{L}},\frac{\epsilon^{\frac{p+2}{4p+3}}}{\kappa^{\frac{3p}{8p+6}}d^{\frac{p}{4p+3}}\sqrt{L}}\right)$ and $N=C_{4}\big{[}\frac{\kappa^{7/6}d^{1/6}}{\epsilon^{1/3}}+\frac{\kappa^{\frac{11p+6}{8p+6}}d^{\frac{p}{4p+3}}}{\epsilon^{\frac{p+2}{4p+3}}}\big{]}$, we have

### v3

@@ 1 Introduction
Sampling from a density $\pi(x)\propto\exp(-F(x))$ over $\mathbb{R}^{d}$ is of fundamental interest in physics, economics, and finance johannes2010mcmc ; von2011bayesian ; kobyzev2020normalizing . Applications in computer science include volume computation vempala2010recent and bandit optimization russo2018tutorial . A popular approach is Langevin Monte Carlo (LMC) which is the Euler-Maruyama discretization of the continuous time Itô Stochastic Differential Equation (SDE) called (overdamped/underdamped) Langevin Dynamics. The convergence of LMC has been extensively studied in the literature durmus2018analysislangevinmontecarlo ; vempala2022rapidconvergenceunadjustedlangevin ; erdogdu2021convergencelangevinmontecarlo ; cheng2017convergencelangevinmcmckldivergence ; cheng2018-underdamped-lmc ; dalalyan2020sampling ; altschuler2025shiftedcompositionivunderdamped under various assumptions on the target density $\pi,$ such as log-concavity and isoperimetry.
@@
Strong $L^{2}$ error is the $L^{2}$ distance between the continuous time Itô SDE solution at time $T$ and the sampling algorithm output whenever they are driven by same Brownian motion. However, sampling algorithm guarantees generally consider ‘weak’ notions of distance such as total variation distance, Wasserstein distance, or the KL divergence between the law of algorithm output and the target. In particular, Wasserstein-2 distance bounds can consider the $L^{2}$ distance between algorithm output and the continuous time SDE driven by different but arbitrarily coupled Brownian motions.
@@ 2 Notation and problem setup
Let $\|\cdot\|$ denote the standard Euclidean norm over $\mathbb{R}^{d}$ for some $d$ and $\mathbf{I}_{d}$ denote the $d\times d$ identity matrix. $x\lesssim y$ means there exists a universal constant $C>0$ such that $x\leq Cy$. Let $\mathsf{Law}(X)$ denote the law of the random variable $X$. Given a probability measure $\mu$, $X\sim\mu$ means $\mu=\mathsf{Law}(X)$. $O_{a}(),\Omega_{a}(),\Theta_{a}()$ are same as $O,\Omega,\Theta$, except that they hide the dependence on $a$. Given two probability measures $\mu$ and $\nu$, we let $\text{KL}(\mu||\nu)$ denote the KL divergence and $\text{TV}(\mu,\nu)$ denote the total variation distance between them.
@@
We define the condition number $\kappa:=\frac{L}{\alpha}$. Our notion of complexity is the number of gradient calls of $F$, in terms of the problem parameters $\kappa,d$ and $\epsilon.$ The remainder of this section closely follows the discussion in kandasamy2024poissonmidpointmethodlangevin , since we have the same problem setup.
2.1 Langevin Monte Carlo
@@
Here $Z_{t}$ are independent standard Gaussians, $b(x,t):\mathbb{R}^{d}\times\mathbb{R}^{+}\to\mathbb{R}^{d}$ is a drift function; $A_{\eta},$ $G_{\eta}$ and $\Gamma_{\eta}$ are $d\times d$ matrices. We focus on two specific cases of the above.
Overdamped LMC: Here $A_{\eta}=\mathbf{I}_{d}$, $G_{\eta}=\eta\mathbf{I}_{d}$, $\Gamma_{\eta}=\sqrt{2\eta}\mathbf{I}_{d},$ and $b(x,t)=-\nabla F(x)$ for some potential $F.$ This is the Euler-Maruyama discretization of Overdamped Langevin dynamics:
@@
$dU_{t}=V_{t}dt,\text{ }dV_{t}=-\gamma V_{t}-\nabla F(U_{t})dt+\sqrt{2}dB_{t}.$ (3)
[…]
@@ 2.2 The Poisson midpoint method
Let $K\in\mathbb{N}$. Let $X_{t}$ denote the dynamics in (2) with step size $\eta/K$. The Poisson midpoint method uses a randomized algorithm to approximate $K$ steps of these dynamics with just 1 step of size $\eta$; using 2 evaluations of the drift $b()$ in expectation. Denoting the iterates of the Poisson midpoint method as $\tilde{X}_{tK}$, this means $\tilde{X}_{tK}$ approximates $X_{tK}$. The work of kandasamy2024poissonmidpointmethodlangevin gives an efficient implementation of PLMC, with 2 evaluations of the drift $b()$ per $K$ steps to obtain $\tilde{X}_{(t+1)K}$ directly from $\tilde{X}_{tK}$. Note that neither the oracle nor computational complexity scale with the batch-size $K$. We consider the ‘fictitious’ interpolation $\hat{\tilde{X}}_{tK+i}$ (which approximates $X_{tK+i}$) to enable mathematical analysis. The iterates run in batches of size $K,$ denoted ${\tilde{X}}_{tK+i}$ with $0\leq i<K$; and the initialization is $\tilde{X}_{0}=X_{0}.$ Let $Y_{tK+i}$ denote independent standard Gaussians. Given $\tilde{X}_{tK}$, the 0-th iterate of a batch, we define the cheap interpolation
@@
We consider this approximation in the two specific cases of Langevin Dynamics to obtain Poisson Langevin Monte Carlo (PLMC).
Overdamped PLMC: Let $Z_{tK+i}$ be independent standard Gaussians, and $X_{tK+i}$ denote the iterates of Langevin Monte Carlo, with step-size $\eta/K.$
@@
Underdamped PLMC: Let $Z_{tK+i}$ be independent standard Gaussians in $\mathbb{R}^{2d},$ and $X_{tK+i}=[U_{tK+i},V_{tK+i}]^{T}$ denote the iterates of underdamped Langevin Monte Carlo with step-size $\eta/K$.
@@ 2.3 Prior work
LMC (durmus2018analysislangevinmontecarlo ) Strongly Log-Concave $W_{2}^{2}\leq\frac{\epsilon^{2}}{\alpha}$ $\frac{\kappa d}{\epsilon^{2}}$
RLMC (shen2019randomizedmidpointmethodlogconcave ; yu2023langevinmontecarlostrongly ) Strongly Log-Concave $W_{2}^{2}\leq\frac{\epsilon^{2}}{\alpha}$ $\frac{\kappa\sqrt{d}}{\epsilon}+\frac{\kappa^{4/3}d^{1/3}}{\epsilon^{2/3}}$
RLMC (altschuler2024shiftedcompositioniiilocal ) Strongly Log-Concave $\mathsf{KL}\leq\epsilon^{2}$ $\frac{\kappa\sqrt{d}}{\epsilon}$
@@
PLMC (Ours) Strongly Log-Concave $W_{2}^{2}\leq\frac{\epsilon^{2}}{\alpha}$ $\frac{\kappa^{4/3}d^{1/3}+\kappa d^{2/3}}{\epsilon^{2/3}}$
@@
LMC (dalalyan2020sampling ) Strongly Log-Concave $W_{2}^{2}\leq\frac{\epsilon^{2}}{\alpha}$ $\frac{\kappa^{3/2}\sqrt{d}}{\epsilon}$
RLMC (shen2019randomizedmidpointmethodlogconcave ; yu2023langevinmontecarlostrongly ) Strongly Log-Concave $W_{2}^{2}\leq\frac{\epsilon^{2}}{\alpha}$ $\frac{\kappa d^{1/3}}{\epsilon^{2/3}}+\frac{\kappa^{7/6}d^{1/6}}{\epsilon^{1/3}}$
@@
PLMC (Ours) Strongly Log-Concave $W_{2}^{2}\leq\frac{\epsilon^{2}}{\alpha}$ $\frac{\kappa^{7/6}d^{1/3}}{\epsilon^{1/3}}+\frac{\kappa^{\frac{11p+6}{8p+6}}d^{\frac{3p+2}{8p+6}}}{\epsilon^{\frac{p+2}{4p+3}}}$
Recent works have focused on the rigorous theoretical analysis of classical and popular MCMC algorithms to establish complexity bounds and theoretical limits. The prototypical case of Overdamped LMC has been studied when the target $\pi$ is strongly log-concave and more generally when $\pi$ satisfies isoperimetric inequalities dalalyan2017theoretical ; durmus2017unadjusted ; durmus2018analysislangevinmontecarlo ; vempala2022rapidconvergenceunadjustedlangevin ; erdogdu2021-tail-growth ; mou2022improved ; sinho-non-log-concave-lmc . Underdamped LMC has been considered as a faster alternative. This case too has been well studied when $\pi$ is strongly log-concave and when $\pi$ satisfies isoperimetric inequalities cheng2018-underdamped-lmc ; dalalyan2020sampling ; ganesh2020faster ; ma2021there ; zhang2023improveddiscretizationanalysisunderdamped ; altschuler2025shiftedcompositionivunderdamped
@@
Note that the literature on MCMC considers 3 notions of convergence: KL-divergence, TV distance and 2-Wasserstein distance. In the case when $\pi$ is strongly log-concave, the Otto-Villani Theorem otto2000generalization shows that $\mathsf{KL}(\mu||\pi)\leq\epsilon^{2}\implies W_{2}^{2}(\mu,\pi)\lesssim\frac{\epsilon^{2}}{\alpha}$ and the Pinsker’s inequality shows that $\mathsf{KL}(\mu||\pi)\leq\epsilon^{2}\implies\mathsf{TV}(\mu,\pi)\lesssim\epsilon$. The condition of $\pi$ satisfying LSI is more general than strong log-concavity of the target. We refer to Table 1 for a detailed comparison of the results.
@@ 2.4 Our contributions
However, sampling algorithm guarantees consider ‘weak errors’ which are distances between $\mathsf{Law}(U_{T}(\omega))$ and $)\mathsf{Law}(A(F,\omega,\tilde{\omega}$)). In particular, the Wasserstein-2 distance is the infimum of $L^{2}$ errors when $U_{T}$ is driven by $B_{t}(\omega)$ and $A(\cdot)$ queries $B_{t}^{\prime}(\omega)$ over all couplings of distinct Brownian motions $B_{t}(\omega)$ and $B^{\prime}_{t}(\omega)$. Our results show Poisson ULMC queries the oracle $\tilde{O}_{\kappa,d}(\frac{1}{\epsilon^{1/3}})$ times in expectation to achieve $W_{2}^{2}(\mathsf{Law}(A(F,\omega,\tilde{\omega})),\pi)\leq\frac{\epsilon^{2}d}{\alpha}$, a quadratic improvement over RLMC.
@@ 3 Results
Let $\tilde{X}_{tK+i}$ denote the iterates of Overdamped PLMC, and $X_{tK+i}$ the iterates of Overdamped LMC with stepsize $\eta/K$. Assume $\eta L\leq 1/8,$ and Assumption 1. Then there exist absolute constants $c_{1}$ and $c_{2}$ such that
@@
Let $\tilde{X}_{t}$ be the iterates of Overdamped PLMC as in Theorem 1. Assume the conditions from Theorem 1. Let $x^{*}$ be the unique minimizer of $F$, and $\tilde{X}_{0}$ satisfy $\mathbb{E}[F(\tilde{X}_{0})-F(x^{*})]\leq C_{F}\kappa d$ for some $C_{F}>0$. Then there exist constants $C_{1},C_{2}>0$ depending only on $C_{F}$, $\log(\frac{W_{2}(U_{0},\pi)\sqrt{\alpha}}{\epsilon\sqrt{d}})$ and $\log(1/\epsilon),$ polynomially, such that if $\eta=C_{1}\min(\frac{\alpha^{1/3}\epsilon^{2/3}}{L^{4/3}},\frac{\epsilon^{2/3}}{d^{1/3}L}),$ $K\asymp\max(\frac{\eta L}{\epsilon^{2}},1)$ and $N=C_{2}\big{[}\frac{\kappa^{4/3}+\kappa d^{1/3}}{\epsilon^{2/3}}\big{]}$. Then,
@@
The complexity bound for Overdamped LMC (durmus2018analysislangevinmontecarlo ) is $\tilde{O}(\kappa/\epsilon^{2})$ gradient calls, and that of Overdamped RLMC (yu2023langevinmontecarlostrongly ) is $\tilde{O}(\frac{\kappa}{\epsilon}+\frac{\kappa^{4/3}}{\epsilon^{2/3}})$ gradient calls. To our knowledge, our method is thus the best known discretization of overdamped Langevin dynamics, in terms of $\epsilon$ dependence. Note that our assumption on the initialization is very mild - $F$ can be optimized quickly under Assumption 1.
@@
Let $\tilde{X}_{tK+i}$ denote the iterates of Underdamped PLMC, and $X_{tK+i}$ denote the iterates of Underdamped LMC with step-size $\eta/K$. Let $M=\begin{bmatrix}\mathbf{I}_{d}&0\\ \mathbf{I}_{d}&\frac{2}{\gamma}\mathbf{I}_{d}\end{bmatrix}$, and $p\geq 0$ be an integer. Assume $\eta/K\lesssim\frac{1}{\kappa\sqrt{L}},$ the conditions of Theorem 6 , and Assumption 1. Assume $\gamma\eta<c_{0}$ for sufficiently small $c_{0}>0,$ depending only on $p.$ Then with $\gamma=c\sqrt{L}$ for some $c\geq 2$, we have
@@
[…]
Where $P$ is a polynomial containing high powers of $\eta$, and $\Psi$ is defined as $\Psi_{s}:=\tilde{U}_{sK}+\frac{1}{\gamma}\tilde{V}_{sK}.$ The complete bound is explicated in Section 6.1, for the sake of clarity.
The bound in Theorem 3 holds for any choice of nonnegative integer $p$. The presence of $p$ is due to the manner in which we bound a certain low probability event - see the proof of Proposition 3. Similar to Corollary 1, the following Corollary (proved in Section F) establishes complexity bounds.
@@
[…]